OpenStudy (narissa):

Fill in the blanks with the reasons for the following proof; prove that -(x+y)+x=-y

OpenStudy (narissa):

Theres a chart on the left says statements on the right says rea under statement it says number 1) -(x+y)+x=-x+(-y)+x and it asking for reason

OpenStudy (narissa):

theres 2), 3), 4) i need to fill in

OpenStudy (narissa):

@phi

OpenStudy (narissa):

2)it says =-x+x+(-y) 3)=0+(-y) and number 4)=-y

OpenStudy (narissa):

@freckles

OpenStudy (freckles):

do you know any properties/laws? such as distributive law, commutative property, associative property, additive inverse property....

OpenStudy (narissa):

i think the first one is commutative

OpenStudy (freckles):

commutative property says it is okay to change the order of terms... so are you saying in the first step they did something like write a+b as b+a

OpenStudy (narissa):

no

OpenStudy (narissa):

what is the one in changing parenthesis

OpenStudy (freckles):

can you state what the distributive law says?

OpenStudy (narissa):

ok then thats the one it dont matter where the paranthesis are

OpenStudy (narissa):

2) i just need to change signs right?

OpenStudy (freckles):

change signs?

OpenStudy (narissa):

1) is distributive right?

OpenStudy (freckles):

well what do you think the distributive law says?

OpenStudy (narissa):

u distribute stuff

OpenStudy (freckles):

a(b+c)=?

OpenStudy (narissa):

ab+ac right?

OpenStudy (narissa):

yes

OpenStudy (freckles):

yes a(b+c)=ab+ac is the distributive law isn't that the same thing that happened here: -1(x+y)=-1x+(-1y) or you can also write as -(x+y)=-x+(-y)

OpenStudy (narissa):

k thanks what about 2 i have to finish filling in number and put reason i think i change the signs and its additive inverse property

OpenStudy (freckles):

change the signs? what do you mean by that?

OpenStudy (narissa):

2) shows _____________=-x+x+(-y)

OpenStudy (freckles):

-x+(-y)+x look what happened to the terms; look at the locations of the terms; -x+x+(-y)

OpenStudy (narissa):

ok they move them around

OpenStudy (freckles):

what property says it is okay to change the order of the terms

OpenStudy (narissa):

what numbers do i fill in though

OpenStudy (freckles):

numbers?

OpenStudy (freckles):

I don't know what numbers you mean

OpenStudy (freckles):

I thought we are just giving the reasons

OpenStudy (narissa):

communitive

OpenStudy (freckles):

right commutative property

OpenStudy (narissa):

i have to give the reasons the first one they filled in the numbers the next 3 i have to fill in the first part to make it = what it says

OpenStudy (narissa):

the first one is done for me says -(x+y)+x=-x+(-y)+x

OpenStudy (narissa):

The second one shows _______________=-x+x+(-y)

OpenStudy (narissa):

third one ____________________=0+(-y)

OpenStudy (narissa):

fourth one _____________________=-y

OpenStudy (narissa):

i have to fill in the first part of each

OpenStudy (narissa):

do u see now?

OpenStudy (narissa):

is it hard because it is foir me?

OpenStudy (narissa):

do u know how to do it?

OpenStudy (narissa):

@FibonacciChick666

OpenStudy (fibonaccichick666):

Didn't we do this already?

OpenStudy (narissa):

its a different one

OpenStudy (narissa):

for the second one do i move the parethensis?

OpenStudy (fibonaccichick666):

http://openstudy.com/study#/updates/56270c1ae4b09cb8ec242ea6

OpenStudy (fibonaccichick666):

-(x+y)+x=-y

OpenStudy (fibonaccichick666):

so what is your first step and why?

OpenStudy (narissa):

im confused

OpenStudy (fibonaccichick666):

ok, where and why?

OpenStudy (narissa):

they moved the parenthesis

OpenStudy (narissa):

did we do this already i cant find it?

OpenStudy (fibonaccichick666):

oh, I did it with jinxhead

OpenStudy (narissa):

ok

OpenStudy (fibonaccichick666):

alright. Let's start with the crappy chart

OpenStudy (freckles):

@FibonacciChick666 he said something about typing in numbers also ...do you have any idea what he means?

OpenStudy (freckles):

I thought we were just suppose to give reasons for each of the steps...

OpenStudy (fibonaccichick666):

so, this idiocy is a two column proof that looks like this. The other person who did this had no idea how to fill it in either.... smh 

OpenStudy (fibonaccichick666):

ok, so the whole thought for this is how do I fill in the chart. What does each number mean

OpenStudy (freckles):

@narissa I don't see any place in that table to put numbers... So what numbers are you talking about typing in?

OpenStudy (narissa):

it is like that but 1)-(x-y)+x=-x+(-y)+x i put distributive and number 2) i have to fill in the first part______=-x+x+(-y)

OpenStudy (narissa):

im just confused

OpenStudy (freckles):

I'm pretty sure @FibonacciChick666 is right all you have to do is put the reason for each step

OpenStudy (freckles):

we already talked about distributive law and commutative property

OpenStudy (fibonaccichick666):

ok, so here is how the chart works. #1 on reasons, is given. Always. #2 reason is how you go from step 1 to step 2

OpenStudy (freckles):

what property mentions -x+x=0

OpenStudy (fibonaccichick666):

#3 reason is how you go from step 2 to step 3

OpenStudy (fibonaccichick666):

etc

OpenStudy (fibonaccichick666):

So number two under reasons is distributive property because to go from step one to step two I used the distributive property.

OpenStudy (fibonaccichick666):

can you fill in 3-5?

OpenStudy (narissa):

ok so 1 is distributive 2 is commutative 3 inverserse and 4 is identity

OpenStudy (narissa):

theres only 1-4

OpenStudy (freckles):

that is 2-5

OpenStudy (fibonaccichick666):

no, remember what I said 1 always is? There is no step zero, so we can't have a property

OpenStudy (fibonaccichick666):

but if that is 2-5, then yes